Suppose the price elasticity of demand for heating oil is 0.2 in short run and 0.7 in the long run

a) if the price of heating oil rises from $1.80 to$2.20 per gallon, what happens to the quantity of heating oil demanded in short run? In the long run?

b) Why might this elasticity depend on the time horizon?


1
Expert's answer
2022-03-07T10:43:44-0500

The price elasticity of demand for heating oil in the short run = 0.2.

The price elasticity of demand for heating oil in the long run = 0.7.


PED- represents Price elasticity of demand/.

The midpoint formula of the price elasticity of demand is;


"PED=\frac {\% \Delta Quantity\space demanded}{\% \Delta Price}"


Where;

"\%\Delta Qd= \frac {Q_2-Q_1} {\frac {Q_1+Q_2} {2}}"


"\%\Delta P= \frac {P_2-P_1} {\frac {P_1+P_2}{2}}"


a)

If the price of heating oil rises from $1.80 to $2.20 per gallon then quantity of heating oil demanded in the short run will be:


"PED= \frac{\frac{\%\Delta Qd}{P_2-P_1}} {\frac {P_1+P_2}{2}}"


"0.2= \frac{\frac{\%\Delta Qd}{2.20-1.80}} {\frac {1.80+2.20}{2}}"


"0.2 =\frac {\%\Delta Qd}{ \frac{0.4}{2}}"


"\%\Delta Qd = \frac {0.2\times 0.4}{2}=0.04"


It, therefore, means that the quantity of heating oil demanded will decrease by 4% in the short run.


The impact of an increase in price on the quantity of heating oil demanded, in the long run, will be:


"0.7= \frac{\frac{\%\Delta Qd}{2.20-1.80}} {\frac {1.80+2.20}{2}}"


"0.7 =\frac {\%\Delta Qd}{ \frac{0.4}{2}}"


"\%\Delta Qd = \frac {0.7\times 0.4}{2}=0.14"


Hence, the quantity of heating oil demanded will decrease by 14% in long run.


b)

Elasticity can be said to depend on the time horizon because there is a possibility of substitutes for heating oil in the market. It is likely that people might prefer other sources of heating oil in the future in case of rising in its prices.